OpenStudy (anonymous):

Andrea visits a casino to play a game of American roulette. She bets only on even numbers. What is the probability that she will win? A-6/19 B-8/19 C-9/19 D-1/2 E-3/38

OpenStudy (anonymous):

American roulette has 38 slots, 18 even, 18 odd, and 0 and 00. Win when get one of the 18 evens out of the total of 38 or p=18/38 = 9/19